# Echo Flow Networks

### v1

@@ 1 Introduction
While capturing long-range dependencies, and ideally learning from the entire history of inputs, is essential for accurate prediction, fully leveraging such historical context remains an open challenge due to two critical limitations. First, computational inefficiency: Transformer-based models like PatchTST (Nie et al., 2023) exhibit quadratic complexity in time and memory (Vaswani et al., 2017), making them impractical for ultra-long sequences. Even with recent efficiency improvements (Jia et al., 2024; Lin et al., 2023; Chi, 2024), these models still rely on backpropagation through time, resulting in high training cost and limited scalability. Second, model limitations: many models struggle with vanishing gradients and suffer from bifurcation issues, leaving significant room for performance improvement using long history (Lim & Zohren, 2021; Zhou et al., 2021; Nie, 2023). (Lim & Zohren, 2021; Zhou et al., 2021; Nie, 2023).
@@ 3.1 Experimental Settings
Evaluation Metrics and Datasets: We evaluate performance using Mean Squared Error (MSE) and Mean Absolute Error (MAE) (Hastie et al., 2009), where lower values are better. Benchmarks include representative TSF datasets: four ETT variants (ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al., 2021), Weather and Traffic (Zeng et al., 2022a), Air Quality (AQ) (De Vito et al., 2008), and Daily Website Visitors (DWV) (Nau, 2021). For ETT, Weather, and Traffic, preprocessing follows (Zeng et al., 2023), while AQ and DWV use a 70/10/20 split for training, validation, and testing.
Baselines: We compare EFNs against a range of strong baselines commonly used in TSF, including Transformer-based models such as PatchTST (Nie et al., 2022), Seg-RNN (Lin et al., 2023), MLP-based models like PatchTSMixer (Ekambaram et al., 2023), linear projection models such as DLinear (Zeng et al., 2023), and large pre-trained models like TimeLLM (Jin et al., 2024). The experimental setup follows the standardized protocol in (Zeng et al., 2023) for fair comparison.
@@ 4 Related Work
Deep Learning for TSF Recent TSF models span CNNs (e.g., MICN (Wang et al., 2023), TimesNet (Wu et al., 2023), ModernTCN (Luo & Wang, 2024)) for local patterns; RNNs (e.g., SegRNN (Lin et al., 2023), WITRAN (Jia et al., 2023)) for sequential modeling but prone to gradient issues; and linear models (e.g., FITS (Xu et al., 2024), SparseTSF (Lin et al., 2024b), CycleNet (Lin et al., 2024a)) which are efficient but less expressive. Transformers (e.g., PatchTST (Nie, 2023), TiDE (Huang, 2023), FiLM (Wang, 2023), BasisFormer (Zhou, 2023), iTransformer (Liu et al., 2024), Leddam (Yu et al., 2024)) achieve high accuracy but with high cost. PatchTST (Nie, 2023) and SegRNN (Lin et al., 2023) perform well but struggle with fine-grained or irregular inputs (Nie, 2023; Lin et al., 2023). TimeKAN (Huang et al., 2025) prioritizes efficiency, while TPGN (Jia et al., 2024) risks future leakage (Chi, 2024). EFNs mitigates these by learning full history without backpropagation through time, combining accuracy with scalability.
ESN Efficiency and Stability ESNs avoid vanishing gradients by using fixed recurrent weights (Jaeger, 2001), requiring no backpropagation through time. They offer constant memory and optimization costs ($\mathcal{O}(1)$) via lightweight readouts. Approaches such as spectral radius tuning (Lukoševičius & Jaeger, 2009), regularization (Rodan & Tiňo, 2011), and reservoir optimization (Lu et al., 2017) aim to stabilize ESNs but often need careful tuning. EFNs instead aggregates multiple randomly initialized Group X-ESNs and integrates them via cross-attention—more adaptive than prior parallel reservoirs (Sun et al., 2024; Casanova et al., 2023; Rodan & Tino, 2011). Furthermore, ESNs suffer from initialization sensitivity and limited readout expressiveness. Prior work explores spectral tuning (Jaeger, 2001; Rodan & Tiňo, 2011), nonlinear readouts (Gauthier et al., 2021a), attention (Ma et al., 2022), and graph-based methods (Liao, 2023). EFNs unifies ESNs with a dual-stream architecture and attention-based readout for improved accuracy and adaptability.

### v2

@@ 1 Introduction
While capturing long-range dependencies, and ideally learning from the entire history of inputs, is essential for accurate prediction, fully leveraging such historical context remains an open challenge due to two critical limitations. First, computational inefficiency: Transformer-based models like PatchTST (Nie et al., 2023) exhibit quadratic complexity in time and memory (Vaswani et al., 2017), making them impractical for ultra-long sequences. Even with recent efficiency improvements (Jia et al., 2024; Lin et al., 2023; Chi, 2024), these models still rely on backpropagation through time, resulting in high training cost and limited scalability. Second, model limitations: many models struggle with vanishing gradients and suffer from bifurcation issues, leaving significant room for performance improvement using long history (Lim & Zohren, 2021; Zhou et al., 2021; Nie et al., 2023). (Lim & Zohren, 2021; Zhou et al., 2021; Nie et al., 2023).
@@ 3.1 Experimental Settings
Evaluation Metrics and Datasets: We evaluate performance using Mean Squared Error (MSE) and Mean Absolute Error (MAE) (Hastie et al., 2009), where lower values are better. Benchmarks include representative TSF datasets: four ETT variants (ETTh1, ETTh2, ETTm1, ETTm2) (Zhou et al., 2021), Weather and Traffic (Zeng et al., 2022b), Air Quality (AQ) (De Vito et al., 2008), and Daily Website Visitors (DWV) (Hyndman, R. J.,, ). For ETT, Weather, and Traffic, preprocessing follows (Zeng et al., 2022b), while AQ and DWV use a 70/10/20 split for training, validation, and testing.
Baselines: We compare EFNs against a range of strong baselines commonly used in TSF, including Transformer-based models such as PatchTST (Nie et al., 2023), Seg-RNN (Lin et al., 2023), MLP-based models like PatchTSMixer (Ekambaram et al., 2023), linear projection models such as DLinear (Zeng et al., 2022b), and large pre-trained models like TimeLLM (Jin et al., 2024). The experimental setup follows the standardized protocol in (Zeng et al., 2022b) for fair comparison.
@@ 4 Related Work
Deep Learning for TSF Recent TSF models span CNNs (e.g., MICN (Wang et al., 2023), TimesNet (Wu et al., 2023), ModernTCN (Luo & Wang, 2024)) for local patterns; RNNs (e.g., SegRNN (Lin et al., 2023), WITRAN (Jia et al., 2023)) for sequential modeling but prone to gradient issues; and linear models (e.g., FITS (Xu et al., 2024), SparseTSF (Lin et al., 2024b), CycleNet (Lin et al., 2024a)) which are efficient but less expressive. Transformers (e.g., PatchTST (Nie et al., 2023), TiDE (Huang, 2023), FiLM (Wang, 2023), BasisFormer (Zhou, 2023), iTransformer (Liu et al., 2024), Leddam (Yu et al., 2024)) achieve high accuracy but with high cost. PatchTST (Nie et al., 2023) and SegRNN (Lin et al., 2023) perform well but struggle with fine-grained or irregular inputs (Nie et al., 2023; Lin et al., 2023). TimeKAN (Huang et al., 2025) prioritizes efficiency, while TPGN (Jia et al., 2024) risks future leakage (Chi, 2024). EFNs mitigates these by learning full history without backpropagation through time, combining accuracy with scalability.
ESN Efficiency and Stability ESNs avoid vanishing gradients by using fixed recurrent weights (Jaeger, 2001), requiring no backpropagation through time. They offer constant memory and optimization costs ($\mathcal{O}(1)$) via lightweight readouts. Approaches such as spectral radius tuning (Lukoševičius & Jaeger, 2009), regularization (Rodan & Tiňo, 2011), and reservoir optimization (Lu et al., 2017) aim to stabilize ESNs but often need careful tuning. EFNs instead aggregates multiple randomly initialized Group X-ESNs and integrates them via cross-attention—more adaptive than prior parallel reservoirs (Sun et al., 2024; Casanova et al., 2023; Rodan & Tiňo, 2011). Furthermore, ESNs suffer from initialization sensitivity and limited readout expressiveness. Prior work explores spectral tuning (Jaeger, 2001; Rodan & Tiňo, 2011), nonlinear readouts (Gauthier et al., 2021b), attention (Ma et al., 2022), and graph-based methods (Liao, 2023). EFNs unifies ESNs with a dual-stream architecture and attention-based readout for improved accuracy and adaptability.